Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 78,000 units of product were as follows:

Standard Costs Actual Costs
Direct materials 226,200 lbs. at $5.40 223,900 lbs. at $5.20
Direct labor 19,500 hrs. at $18.30 19,950 hrs. at $18.50
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 20,350 direct
labor hrs.:
Variable cost, $2.90 $55,980 variable cost
Fixed cost, $4.60 $93,610 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Material Price Variance $
Direct Materials Quantity Variance $
Total Direct Materials Cost Variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance $
Direct Labor Time Variance $
Total Direct Labor Cost Variance $

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $
Fixed factory overhead volume variance $
Total factory overhead cost variance $
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
223900 $5.20 $1,164,280.00 223900 $5.40 $1,209,060.00 226200 $5.40 $1,221,480.00
-$44,780.00 F -$12,420.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance -$44,780.00 F
Direct material quantity variance -$12,420.00 F
Direct material cost variance -$57,200.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
19950 $18.50 $369,075.00 19950 $18.30 $365,085.00 19500 $18.30 $356,850.00
$3,990.00 U $8,235.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $3,990.00 U
Direct Labor Efficiency variance $8,235.00 U
Direct labor cost variance $12,225.00 U

Solution c:

Variable factory overhead controllable variance = Actual cost of variable overhead - Standard cost of variable overhead

= $55,980 - (19500*$2.90) = - $570 F

Fixed factory overhead volume variance = Budgeted fixed overhead - Fixed overhead applied

= $93,610 - (19500 * $4.60) = $3,910 U

Total factory overhead cost variance = - $570 F + $3,910 U = $3,340 U